Write down expectation value of r^2.
1
Expert's answer
2020-11-10T07:08:29-0500

The expectation value of r2:


"\big<r^2\big>=\int^\infty_0\psi_{1s}(r)r^2\psi_{1s}(r)4\pi r^2\text{d}r=3,"

or 3a02.